OpenStudy (anonymous):

Write the system x′=e7tx−3ty+5sin(t), y′=7tan(t)y+5x−7cos(t) in the form ddt[xy]=P(t)[xy]+f⃗ (t). Use prime notation for derivatives and write x and x′, etc., instead of x(t), x′(t), or dxdt.

OpenStudy (anonymous):

question not clear, can u please click n attach?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

http://i.imgur.com/TxQQt.jpg

OpenStudy (mr.math):

\[\left(\begin{matrix}x' \\ y'\end{matrix}\right)=\left[\begin{matrix}e^{7t} & 5\\ -3t & 7\tan{t}\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)+\left(\begin{matrix}5\sin{t} \\ -7\cos{t}\end{matrix}\right)\]

OpenStudy (anonymous):

i think that should work,thanks so much

OpenStudy (mr.math):

No problem!

OpenStudy (anonymous):

yeah got the same as mr. Math got

OpenStudy (anonymous):

then thank you too